Find internal rate of return for the project whose cash flows are as flows: 8
Initial investment = Rs.5,00,000
Annual benefit = Rs. 1,50,000
Annual cost = Rs.30,000
Salvage value = Rs. 40,000
MARR = 12% per year
1
Expert's answer
2020-06-11T09:23:24-0400

Computing Internal rate of return using trial and error method;

steps;

  • Compute Net present value of the project using an arbitrary selected discount rate.
  • If the NPV so calculated is positive then try a higher rate and if negative try a lower rate.
  • Continue the process until the NPV is equal to zero.
  • Use linear interpolation to determine the exact rate

Formulas to be applied are as below,

"pvif= {1 \above{2pt} \left(1+r )^{\smash{n}}\right)}"


"pvifa= {1-(1+r )\left(^{\smash{n}}\right) \above{2pt} r}"


"IRR=Rl+{NPVh*(Rh-Rl) \above{2pt} NPVh+NPVl}"

where

r -rate of return

n -period in years

Rl-lower rate of return

Rh-higher rate of return

NPVh-Net present value using higher rate of return

NPVl-Net present value using lower rate of return


We select arbitrary discount rate, say 20.5% Assuming 10 year of usage.

since year one to year nine cash flows are in annuity,we first calculate NPV for the annuity as follows;

"net cash flow=inflows-annual cost"


"150,000-30,000=120,000"


"pva=120,000*{1-(1+0.205 )\left(^{\smash{9}}\right) \above{2pt} 0.205}=476,085"


we then calculate present value for the 10th year,


"net cash flow=inflows-annual cost +salvage value"


"150,000\u221230,000+40,000=160,000"


"pv=160,000\u2217(1+0.205)10)1=24,788"


"NPV=PV cash inflows- PV outflows"


"NPV=476,085+24,788-500,000=874"


NPV using a rate of 20.5% is 874,since it is positive but not large we can try a closer higher rate

remember our target is to have NPV at zero

Lets try 20.6, we apply the same concept as above


"pva=120,000*{1-(1+0.206 )\left(^{\smash{9}}\right) \above{2pt} 0.206}=474,576"


"pv=160,000* {1 \above{2pt} \left(1+0.206 )^{\smash{10}}\right)}=24,576"


"NPV=474,085+24,576-500,000=-848"


NPV using a rate of 20.6% is 848

since we have two rate which are close to zero we can use formula 3 above to calculate the exact IRR


"IRR=20.5+{874*(20.6-20.5) \above{2pt} 874--848} =20.55"


"IRR=20.55\%"